Four angles of a Pentagon are 110° 99° 98° and 121°,calculate the fifth angles
Found 2 solutions by ikleyn, josgarithmetic:
Answer by ikleyn(52824)

.
The sum of interior angles of a pentagon is equal to (5-2)*180 = 3*180 = 540 degrees.


It gives you an equation

110° + 99° + 98° + 121° + x = 540°


for the missed 5-th angle, which implies

x = 540 - (110 + 99 + 98 + 121) = 112 degrees  ANSWER.

Solved.

----------------
For many other solved problem on missed angles see the lesson
    - Solved problems on missed angle of a polygon
in this site.


Answer by josgarithmetic(39621)

The pentagon can be cut into three triangles so the total sum of the measures of the interior angles is 3*180 degrees, or 540 degrees.

Your unknown angle: 540-110-99-98-121